OpenStudy (anonymous):

word problem: I run 7 times as far as I swim. I bike 3 times as far as I run. One week I totaled 232 miles. How far did I run?

OpenStudy (asnaseer):

r = miles run s = miles swam b = miles biked r = 7s b = 3r total = b + r + s = 232 therefore: 3r + r + (r/7) = 232 <--- solve this to find r

OpenStudy (anonymous):

I am supposed to solve using a system of equations. I see how to solve the 2nd equation you wrote, but how can I solve by elimination or substitution?

OpenStudy (asnaseer):

we eliminated b and s to get the last equation.

OpenStudy (asnaseer):

using b = 3r and s = r / 7

OpenStudy (asnaseer):

basically you have 3 equations: 1. r = 7s 2. b = 3r 3. b + r + s = 232

OpenStudy (asnaseer):

we used 1 and 2 to eliminate b and s from the 3rd equation.

OpenStudy (anonymous):

That makes sense! Thank you!!

OpenStudy (asnaseer):

yw